OpenStudy (anonymous):

4. a) Show the formular \(\Large \sum_{k=1}^{n}k^{3} = 1 + 2^{3} + 3^{3} +…+n^{3} = (\frac{n(n+1)}{2})^{2}\) by induction over \(\Large n\geq 1\)

OpenStudy (anonymous):

It is true for n=1. Suppose it is true for n. Prove that is true for n+1

OpenStudy (anonymous):

\[\Large \sum_{k=1}^{n+1}k^{3} = 1 + 2^{3} + 3^{3} +…+n^{3} = \\(\frac{n(n+1)}{2})^{2}+(n+1)^3 =(n+1)^2\left ( \frac{ n^2} {2^2} +(n+1)\right)=\\ (n+1)^2\left ( \frac{ n^2+4n+4} {2^2}\right)=\\ (n+1)^2\left ( \frac{(n+2)^2} {2^2}\right)= \] We are done.

OpenStudy (anonymous):

thank you very much Mr Elias

OpenStudy (anonymous):

yw